ashari75:

Below are the steps to solve an equation: Step 1: |x − 5| + 2 = 5 Step 2: |x − 5| = 5 − 2 Step 3: |x − 5| = 3 Which of the following is a correct next step to solve the equation?

Kaihalojr:

what do you think the next step is

ashari75:

i think the next is to remove the brakits

Kaihalojr:

those just show absolute value

Kaihalojr:

so basically 5 and x can't be negative values

ashari75:

so the answer will be x-5=-3

Kaihalojr:

how did you get that?

ashari75:

because it was multiple choice question and look at my answer choices and compared it to what you just told me

Kaihalojr:

I don't see any answer choices

Kaihalojr:

is it an image file if so use the attach file button to post the picture

ashari75:

ik cause i didnt put it on there i just copied and paste the question

Kaihalojr:

Can you please show me the answer choices?

ashari75:

A.x+5=-3 B.-x-5=3 C.x-5=-3 D.x+5=3

ashari75:

thanks u very very helpful

Kaihalojr:

I try

ashari75:

thanks for the medal

Kaihalojr:

honestly lots of people would have been better at helping than me

Kaihalojr:

an example is Vocaloid she's very smart at least compared to me

ashari75:

thx

Vocaloid:

If you have an equation in the format |x - A| = B You can split the equation into x - A = B and x - A = -B Try to apply this logic to your problem to see which equation is one of the correct solutions